OpenStudy (anonymous):

calculate the no. of moles : (a) 392 g of H2SO4. (b) 44.8 L of CO2. (c) 9 g Al. (d) 25 g CaCo3.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Can you try it once??

OpenStudy (anonymous):

let me c

OpenStudy (anonymous):

Find molecular mass of H2SO4 then use molar mass concept..

OpenStudy (anonymous):

no of moles in 392 g of H2SO4 = x

OpenStudy (anonymous):

am not getting !

OpenStudy (anonymous):

you better teach !

OpenStudy (anonymous):

See: Find molecular mass of H2SO4..

OpenStudy (anonymous):

98

OpenStudy (anonymous):

we got that in the first problem :)

OpenStudy (anonymous):

1 mole = 98 g Right using molar mass..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

98 g = 1 mole Find it for 392 g..

OpenStudy (anonymous):

392 * 1 / 98 = 4 moles correct ?

OpenStudy (anonymous):

98*4 = 392 I think..

OpenStudy (anonymous):

Yep you are right..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

c and d are same so lets leave it

OpenStudy (anonymous):

i'll do it on my own :)

OpenStudy (anonymous):

For b: 22.4 L = 1 mole Find it for 44.8 L.

OpenStudy (anonymous):

how to do b ?

OpenStudy (anonymous):

1 * 44.8 /22.4 = 2 moles

OpenStudy (anonymous):

Now do here all the three left and I am going to have bath.. While teaching you I forgot to take bath today.. Yes 2 moles is right @ashna Now try other two till then I am having bath..

OpenStudy (anonymous):

okay teacher :P

OpenStudy (anonymous):

I am here only.. So I will check them when I will come.. So don't worry...

OpenStudy (anonymous):

thank you teacher :)

OpenStudy (anonymous):